# EHR-R1: A Reasoning-Enhanced Foundational Language Model for Electronic Health Record Analysis

### v1

@@ 4.3 Model Training
For our experiments, we adopt Qwen3-1.7B, Qwen3-8B, and Qwen2.5-72B [34] as our base models. The same three stage training strategy us applied to all three models. All the training are conduct based on the code of Transformers [^4] https://github.com/huggingface/transformers and TRL [^5] https://github.com/huggingface/trl framework. For GRPO stage, The weighted hyperparameters of rewards $\lambda_{\mathrm{fmt}}$ and $\lambda_{\mathrm{acc}}$ are both set to 1 and the number of rollout is 8 per sample. More detail training hyperparameters are shown in Table 1.
@@ 4.4.1 Benchmarks
EHRSHOT, a public dataset collected by Stanford University, comprises electronic health records from 7,000 patients. Its publicly available portion includes 14 risk-prediction tasks categorized into three subtypes: Operational Outcomes, Anticipating Lab Test Results, and Assignment of New Diagnoses. EHRSHOT is originally designed for traditional machine learning models, presents a challenge for Large Language Models (LLMs) due to its direct use of various medical codes, which are inherently difficult for LLMs to interpret. To address this, we’ve reformatted the EHRSHOT data into a Markdown-like free-text, similar to EHR-Bench, enabling direct testing with LLMs. Specifically, we utilized the descriptive mapping lexicon provided within EHRSHOT to convert item codes into natural language text. However, certain codes, such as CPT4 and ICD10PCS, could not be directly transformed. Following previous work on evaluating LLMs with EHRSHOT [21], we manually added extra mapping files to resolve these inconsistencies [^6] CPT4 mapping: https://gist.github.com/lieldulev/439793dc3c5a6613b661c33d71fdd185
@@ 4.4.2 Metrics
For risk-prediction tasks, we follow prior work [19] and adopt the area under the receiver operating characteristic curve (AUROC) as the evaluation metric. AUROC measures a model’s ability to distinguish between positive and negative outcomes across all decision thresholds, and can be written as
@@ 4.4.3 Baselines
Developed by Alibaba Cloud, the Qwen2.5/3 series [34] offers versatile, general-purpose multilingual models with a unique ‘thinking’ and ‘non-thinking’ hybrid reasoning engine and Mixture-of-Experts (MoE) architecture for efficient inference. Trained on 36 trillion tokens across over 100 languages, Qwen3 excels in general reasoning, coding, and agent capabilities. While not specifically medical, its robust multilingual support and advanced reasoning could serve as a powerful general foundation for processing diverse medical literature or assisting in global health information dissemination.
@@
Created by Google, MedGemma [37] is a collection of Gemma 3 variants specifically optimized for medical text and image comprehension. Its key innovation lies in its multimodal capability, seamlessly integrating natural language processing with computer vision to analyze medical images (e.g., X-rays, histopathology) alongside textual patient data. Trained on extensive de-identified medical datasets, MedGemma demonstrates superior performance on clinical benchmarks like MedQA and is designed to accelerate healthcare AI applications while supporting local deployment for data privacy.
@@
The Llama3 Series [32], developed by Meta, comprises a family of powerful open-source general-purpose LLMs ranging from 8 billion to 405 billion parameters, with evolving multimodal capabilities in Llama 3.2. Trained on over 15 trillion tokens, these models excel in broad applications, including complex reasoning, coding, and multilingual understanding. While not inherently specialized for medicine, the Llama3 architecture serves as a robust foundation, with fine-tuned variants like Me-LLaMA demonstrating superior performance in various medical question-answering and summarization tasks, highlighting its adaptability for domain-specific healthcare solutions.
@@
Developed by Saama AI Labs, OpenBioLLM [38] is an 8-billion-parameter open-source language model specifically designed for the biomedical field, built upon the Llama 3 architecture. It is meticulously fine-tuned on high-quality biomedical data, enabling it to accurately understand and generate text for tasks like clinical entity recognition, medical question answering, and patient data de-identification. OpenBioLLM consistently outperforms larger general-purpose models on diverse biomedical benchmarks, making it a highly effective and transparent choice for healthcare and life sciences applications.
@@
Baichuan-M2 [39], from Baichuan Intelligent, stands out as the first open-source LLM specifically optimized for medical scenarios, trained "from scratch" on an unprecedented 20 trillion tokens of high-quality medical and general data. This dedicated training, coupled with a multi-stage curriculum learning approach, enables it to achieve deep medical expertise, outperforming models five times its size in clinical practice benchmarks while maintaining strong general capabilities. Its specialized focus makes it highly relevant for tasks such as clinical decision support, medical education, and rare disease diagnosis.
@@
OpenAI’s GPT-4o (omni) [27] is a flagship multimodal model designed for natural human-computer interaction, accepting and generating any combination of text, audio, image, and video through a single, end-to-end neural network . It offers real-time responsiveness and enhanced tokenization for non-English languages, making it faster and more cost-effective. While a general-purpose model, its advanced multimodal understanding and reasoning capabilities could be applied to complex medical cases involving diverse data types, from patient consultations (audio/text) to diagnostic images.
@@
Developed by DeepSeek, R1 [35] is an open-source AI model released in January 2025, primarily focused on pushing the limits of reinforcement learning as a post-training technique for complex reasoning. Utilizing a Mixture-of-Experts (MoE) architecture, it achieves high accuracy in mathematics, programming, and general reasoning tasks at a significantly lower operating cost. Notably, DeepSeek R1 has shown strong performance on medical benchmarks like MedQA [45], demonstrating its potential for efficient and accurate reasoning in healthcare contexts.
@@
The GPT-OSS (Open-Source Series) [36] is OpenAI’s first open-source GPT-class model since GPT-2, released under an Apache 2.0 license to provide powerful reasoning and agentic capabilities to the open community. The models, gpt-oss-120b and gpt-oss-20b, utilize a Mixture-of-Experts (MoE) architecture for efficient inference and support a long 128k token context window. While a general-purpose reasoning model, its performance on complex, health-related queries is noteworthy; the gpt-oss-120b model achieved a score of 30% on the HealthBench benchmark, outperforming OpenAI’s own o4-mini and o3-mini models on this specific task. This demonstrates its potential as a strong, customizable foundation for a variety of applications, including those with intricate medical reasoning requirements.

### v2

@@ 4.3 Model Training
For our experiments, we adopt Qwen3-1.7B, Qwen3-8B, and Qwen2.5-72B [yang2025qwen3] as our base models. The same three stage training strategy us applied to all three models. All the training are conduct based on the code of Transformers [^4] https://github.com/huggingface/transformers and TRL [^5] https://github.com/huggingface/trl framework. For GRPO stage, The weighted hyperparameters of rewards $\lambda_{\mathrm{fmt}}$ and $\lambda_{\mathrm{acc}}$ are both set to 1 and the number of rollout is 8 per sample. More detail training hyperparameters are shown in Table 1.
@@ 4.4.1 Benchmarks
EHRSHOT, a public dataset collected by Stanford University, comprises electronic health records from 7,000 patients. Its publicly available portion includes 14 risk-prediction tasks categorized into three subtypes: Operational Outcomes, Anticipating Lab Test Results, and Assignment of New Diagnoses. EHRSHOT is originally designed for traditional machine learning models, presents a challenge for Large Language Models (LLMs) due to its direct use of various medical codes, which are inherently difficult for LLMs to interpret. To address this, we’ve reformatted the EHRSHOT data into a Markdown-like free-text, similar to EHR-Bench, enabling direct testing with LLMs. Specifically, we utilized the descriptive mapping lexicon provided within EHRSHOT to convert item codes into natural language text. However, certain codes, such as CPT4 and ICD10PCS, could not be directly transformed. Following previous work on evaluating LLMs with EHRSHOT [hegselmann2025large], we manually added extra mapping files to resolve these inconsistencies [^6] CPT4 mapping: https://gist.github.com/lieldulev/439793dc3c5a6613b661c33d71fdd185
@@ 4.4.2 Metrics
For risk-prediction tasks, we follow prior work [NEURIPS2023_d42db1f7] and adopt the area under the receiver operating characteristic curve (AUROC) as the evaluation metric. AUROC measures a model’s ability to distinguish between positive and negative outcomes across all decision thresholds, and can be written as
@@ 4.4.3 Baselines
Developed by Alibaba Cloud, the Qwen2.5/3 series [yang2025qwen3] offers versatile, general-purpose multilingual models with a unique ‘thinking’ and ‘non-thinking’ hybrid reasoning engine and Mixture-of-Experts (MoE) architecture for efficient inference. Trained on 36 trillion tokens across over 100 languages, Qwen3 excels in general reasoning, coding, and agent capabilities. While not specifically medical, its robust multilingual support and advanced reasoning could serve as a powerful general foundation for processing diverse medical literature or assisting in global health information dissemination.
@@
Created by Google, MedGemma [sellergren2025medgemma] is a collection of Gemma 3 variants specifically optimized for medical text and image comprehension. Its key innovation lies in its multimodal capability, seamlessly integrating natural language processing with computer vision to analyze medical images (e.g., X-rays, histopathology) alongside textual patient data. Trained on extensive de-identified medical datasets, MedGemma demonstrates superior performance on clinical benchmarks like MedQA and is designed to accelerate healthcare AI applications while supporting local deployment for data privacy.
@@
The Llama3 Series [DBLP:journals/corr/abs-2407-21783], developed by Meta, comprises a family of powerful open-source general-purpose LLMs ranging from 8 billion to 405 billion parameters, with evolving multimodal capabilities in Llama 3.2. Trained on over 15 trillion tokens, these models excel in broad applications, including complex reasoning, coding, and multilingual understanding. While not inherently specialized for medicine, the Llama3 architecture serves as a robust foundation, with fine-tuned variants like Me-LLaMA demonstrating superior performance in various medical question-answering and summarization tasks, highlighting its adaptability for domain-specific healthcare solutions.
@@
Developed by Saama AI Labs, OpenBioLLM [dorfner2024biomedical] is an 8-billion-parameter open-source language model specifically designed for the biomedical field, built upon the Llama 3 architecture. It is meticulously fine-tuned on high-quality biomedical data, enabling it to accurately understand and generate text for tasks like clinical entity recognition, medical question answering, and patient data de-identification. OpenBioLLM consistently outperforms larger general-purpose models on diverse biomedical benchmarks, making it a highly effective and transparent choice for healthcare and life sciences applications.
@@
Baichuan-M2 [wang2025baichuan], from Baichuan Intelligent, stands out as the first open-source LLM specifically optimized for medical scenarios, trained "from scratch" on an unprecedented 20 trillion tokens of high-quality medical and general data. This dedicated training, coupled with a multi-stage curriculum learning approach, enables it to achieve deep medical expertise, outperforming models five times its size in clinical practice benchmarks while maintaining strong general capabilities. Its specialized focus makes it highly relevant for tasks such as clinical decision support, medical education, and rare disease diagnosis.
@@
OpenAI’s GPT-4o (omni) [hurst2024gpt] is a flagship multimodal model designed for natural human-computer interaction, accepting and generating any combination of text, audio, image, and video through a single, end-to-end neural network . It offers real-time responsiveness and enhanced tokenization for non-English languages, making it faster and more cost-effective. While a general-purpose model, its advanced multimodal understanding and reasoning capabilities could be applied to complex medical cases involving diverse data types, from patient consultations (audio/text) to diagnostic images.
@@
Developed by DeepSeek, R1 [DBLP:journals/corr/abs-2501-12948] is an open-source AI model released in January 2025, primarily focused on pushing the limits of reinforcement learning as a post-training technique for complex reasoning. Utilizing a Mixture-of-Experts (MoE) architecture, it achieves high accuracy in mathematics, programming, and general reasoning tasks at a significantly lower operating cost. Notably, DeepSeek R1 has shown strong performance on medical benchmarks like MedQA [jin2021disease], demonstrating its potential for efficient and accurate reasoning in healthcare contexts.
@@
The GPT-OSS (Open-Source Series) [gptoss] is OpenAI’s first open-source GPT-class model since GPT-2, released under an Apache 2.0 license to provide powerful reasoning and agentic capabilities to the open community. The models, gpt-oss-120b and gpt-oss-20b, utilize a Mixture-of-Experts (MoE) architecture for efficient inference and support a long 128k token context window. While a general-purpose reasoning model, its performance on complex, health-related queries is noteworthy; the gpt-oss-120b model achieved a score of 30% on the HealthBench benchmark, outperforming OpenAI’s own o4-mini and o3-mini models on this specific task. This demonstrates its potential as a strong, customizable foundation for a variety of applications, including those with intricate medical reasoning requirements.
@@ 6 Code Availability
7 Acknowledgments
This work is supported by the National Key R&D Program of China (No. 2022ZD0160702), and the Scientific Research Innovation Capability Support Project for Young Faculty (ZY-GXQNJSKYCXNLZCXM-I22).
8 Author Contributions
All listed authors clearly meet the ICMJE 4 criteria. Y.L., C.W., and J.L. contribute equally to this work. Y.W. and W.X. are the corresponding authors. Specifically, Y.L., C.W., J.L., S.J., P.Q., H.W., Y.Y., S.Z., J.W., Q.F., J.G., Y.Z., Y.W., Y.W., and W.X. all make contributions to the conception or design of the work, and Y.L., C.W., J.L. further perform acquisition, analysis, or interpretation of data for the work. In writing, Y.L., C.W., J.L. draft the work. S.J., P.Q., H.W., Y.Y., S.Z., J.W., Q.F., J.G., Y.Z., Y.W., Y.W., and W.X. review it critically for important intellectual content. All authors approve of the version to be published and agree to be accountable for all aspects of the work to ensure that questions related to the accuracy or integrity of any part of the work are appropriately investigated and resolved.
9 Supplementary
9.1 Details of Experiments
Supplementary Table 1: Accurate results of decision making tasks in EHR-Bench. The performance are measured with the metric F1 score and shown in in the format ‘mean $\pm$ std’.
Tasks Llama3.3-70B GPT-4o Qwen2.5-72B OpenBioLLM-70B Baichuan-M2-32B Medgemma-27B GPT-OSS-120B DeepSeek-R1 Qwen3-235B EHR-R1-72B
[…]
Supplementary Table 2: Accurate results of ablation experiments. The table presents the performance of different model configurations across various decision-making tasks. We evaluate our method on base models of different sizes, including Qwen3-1.7B, Qwen3-8B, Qwen2.5-72B, and Qwen3-235B. The performances are measured with the metric F1 score and shown in the format ‘mean $\pm$ std’, with the average performance across all tasks provided in the final column.
Decision Making Tasks
Methods Transfer Service Procedure Test&Exam Diagnose Treatment ICU Event Average
Qwen3-1.7B
[…]
Qwen3-8B
[…]
Qwen2.5-72B
[…]
9.2 Details of EHR-Bench
Supplementary Table 3: Details of task information in EHR-Bench. This table presents a comprehensive overview of the 42 tasks, categorized into 2 main Task Types and further divided into 12 SubTypes. The Target Event column specifies the type of event associated with the target label for each task. For decision making tasks, the Item Name indicates the column name in MIMIC-IV where the target label is located. It’s important to note that Risk Prediction tasks do not have an Item Name as their target labels are not specific entities. Columns marked with ‘*’ represent additional columns created through manual mapping. The Candidates column shows the size of the target candidate set for each task.
Task Type SubType Task Target Event Item Name Metric Candidates
Decision Making Reassignment Admissions admissions admission_type F1 8
Transfer transfer eventtype F1 39
Test & Exam OMR omr result_name F1 11
Labevents labevents item_name F1 698
Microbiologyevents microbiologyevents test_name F1 165
Radiology radiology exam_name F1 961
Diagnoses Diagnose ICD diagnose_icd diagnoses F1 24467
Diagnose CCS diagnose_icd CCT Type* F1 279
Diagnosis ICD diagnosis_icd icd_title F1 13171
Diagnosis CCS diagnosis_icd CCS Type* F1 271
Procedures Procedures ICD procedures_icd procedures F1 11098
Procedures CCS procedures_icd CCS Type* F1 230
Services Services services curr_service F1 18
POE poe order_type F1 15
Treatments EMAR emr medication F1 4153
Prescriptions prescriptions drug F1 9233
Prescriptions ATC prescriptions ATC Type* F1 913
Medrecon medrecon name F1 18641
Medrecon ATC medrecon ATC Type* F1 899
ICU Event Ingredientevents ingredientevents item_name F1 15
Datetimeevents datetimeevents item_name F1 137
Procedureevents procedureevents item_name F1 138
Inputevents inputevents item_name F1 222
Outputevents outputevents item_name F1 63
Risk Prediction Transfer ED Hospitalization edstays - AUROC 2
ED ICU Tranfer 12hour edstays - AUROC 2
Critical Outcomes ED Critical Outcomes edstays - AUROC 2
Readmission Readmission 30day discharge - AUROC 2
Readmission 60day discharge - AUROC 2
ICU Readmission icustays - AUROC 2
ED Reattendance 3day edstays - AUROC 2
LengthOfStay LengthOfStay 3day admissions - AUROC 2
LengthOfStay 7day admissions - AUROC 2
ICU Stay 7day icustays - AUROC 2
ICU Stay 14day icustays - AUROC 2
Mortality ED Inpatient Mortality edstays - AUROC 2
Inpatient Mortality admissions - AUROC 2
ICU Mortality 1day icustays - AUROC 2
ICU Mortality 2day icustays - AUROC 2
ICU Mortality 3day icustays - AUROC 2
ICU Mortality 7day icustays - AUROC 2
ICU Mortality 14day icustays - AUROC 2
9.3 Prompt Collection
Prompt 1.
Reasoning Chain Evaluation Instruction for Clinicians
Task Description: Evaluation of Clinical Reasoning Chain Validity
I. Background
We have developed a large language model (LLM) designed to extract key information from electronic health records (EHRs) and perform clinical reasoning based on that information. For evaluation purposes, we have hidden the full EHR context, presenting only the critical information extracted by the model and the final diagnostic conclusion. Your task is to assess the relationship and logical consistency between the extracted information and the final conclusion.
II. Data Description
The model’s response consists of three main sections: ##Extraction, ##Reasoning, and ##Final Result. You need to evaluate whether the content in ##Extraction and ##Reasoning adequately supports the ##Final Result and provide your final score in the Score column.
III. Evaluation Objective
Using the provided extracted information, you must determine if it is sufficient and accurate to support the model’s final diagnostic conclusion. Specifically, you need to confirm:
- Whether the key information extracted by the model is relevant to the diagnosis and accurately expressed.
- Whether the final diagnostic conclusion is fully supported by the extracted information, without any missing essential evidence.
Note: The model’s reasoning result is considered correct and does not need to be validated for correctness.
IV. Evaluation Process
Please read the entire reasoning chain, including the extracted information and the diagnostic conclusion. Based on your professional judgment, evaluate the degree of correspondence between the information and the conclusion. The scoring criteria are as follows (five levels):
- 5 points (Very Satisfied): The extracted information is mostly accurate and largely supports the conclusion.
- 4 points (Satisfied): The extracted information supports most of the conclusion. A small part of the conclusion lacks evidence but is reasonably justified.
- 3 points (Average): The extracted information supports only a small part of the conclusion. The final result contains a large portion of content that is not adequately supported.
- 2 points (Poor): The extracted information is related to the final result, but it fails to support the conclusion.
- 1 point (Very Poor): The extracted information is highly irrelevant to the diagnostic conclusion. The conclusion lacks a reasonable basis, and the reasoning process is unreliable.
Scoring Differences Explained:
- The main difference between 5 points and 4 points is whether the information extracted by the model fully supports the final result.
- The difference between 5 points and 4 points is whether the final conclusion is fully supported by the extracted information.
- The difference between 4 points and 3 points is whether a large part or a small part of the final conclusion is supported.
- The difference between 3 points and 2 points is whether at least one conclusion is supported. If so, it’s 3 points; if not, but the information is related, it’s 2 points. - The difference between 2 points and 1 point is whether the extracted and reasoned information is relevant to the final conclusion.
V. Evaluation Examples
Example 1 (5 points, Lab Test)
1. Task Type:
Predict the lab tests the patient should undergo.
2. Model Reasoning Chain:
##Extraction:
Source: Outpatient Record [2178-09-23 11:20:00]
- Key Information: Patient complains of fatigue and pale skin/mucous membranes.
- Analysis: Typical signs of anemia, requiring clarification of anemia type.
Source: Lab Results [2178-09-22 15:45:00]
- Key Information: Hemoglobin 8.2 g/dL (normal 13-17), Creatinine 2.1 mg/dL (normal 0.7-1.3).
- Analysis: Moderate to severe anemia combined with abnormal kidney function suggests possible renal anemia.
Source: Past Medical History [2177-05-12 09:00:00]
- Key Information: History of chronic kidney disease (CKD stage 3), not on dialysis.
- Analysis: Chronic kidney disease often leads to renal anemia. Erythropoietin levels need to be assessed.
Source: Past Surgical History [2176-08-05 10:00:00]
- Key Information: Appendectomy, good recovery.
- Analysis: Not directly related to the current anemia evaluation. This is irrelevant information.
##Reasoning:
Combining "anemia symptoms + significantly low hemoglobin + abnormal kidney function + chronic kidney disease," the following lab tests are recommended to clarify the anemia mechanism:
Serum iron
Ferritin
Transferrin saturation
Erythropoietin levels
##Final Result:
Recommended Tests: Serum iron, Ferritin, Transferrin saturation, Erythropoietin levels.
3. Score Explanation:
Score: 5 points
Reasoning:
The first three pieces of extracted information ("fatigue + pale skin," "Hemoglobin 8.2 + Creatinine 2.1," "chronic kidney disease history") fully support the necessity of performing tests for iron metabolism and erythropoietin levels.
The "past appendectomy" is extraneous information unrelated to the conclusion, but it does not affect the correctness of the conclusion.
All key points of the conclusion are fully justified by the first three essential pieces of information, thus a score of 5 points.
Example 2 (4 points, Medication Prescription)
1. Task Type:
Predict the patient’s medication plan.
2. Model Reasoning Chain:
##Extraction:
Source: Inpatient Record [2180-02-15 08:10:00]
- Key Information: Type 2 diabetes mellitus, fasting blood glucose 12.5 mmol/L, postprandial blood glucose 16.0 mmol/L.
- Analysis: Blood glucose is significantly high; the current oral medication regimen is insufficient.
Source: Lab Results [2180-02-14 07:50:00]
- Key Information: HbA1c 8.5%, Creatinine 60 µmol/L (normal).
- Analysis: High HbA1c indicates the need for improved blood glucose control. Normal kidney function means GLP-1 agonists are a viable option.
Source: Past Medication History [2179-10-10 09:00:00]
- Key Information: Taking oral Metformin 1000 mg twice daily and Gliclazide 80 mg once daily.
- Analysis: The current dual-drug therapy has failed to achieve the target, requiring an upgrade.
##Reasoning:
Poor glycemic control (fasting 12.5, postprandial 16.0, HbA1c 8.5%) → Needs escalation of antidiabetic treatment.
Normal kidney function → GLP-1 agonists are a safe option.
Has a history of hypertension and is taking Amlodipine → Needs to continue antihypertensive therapy.
##Final Result:
Prescription: Add a GLP-1 receptor agonist, continue Amlodipine treatment.
3. Score Explanation:
Score: 4 points
Reasoning:
Support for adding GLP-1: The information "fasting 12.5 mmol/L, HbA1c 8.5%" clearly suggests that the current oral medications are insufficient, and an escalation to a GLP-1 is reasonable. Additionally, "normal kidney function" proves this drug can be used safely. This part of the conclusion aligns perfectly with the extracted information.
Support for continuing Amlodipine: The extracted information "history of hypertension and is taking Amlodipine" indicates the patient is already on antihypertensive treatment. This sub-conclusion is also directly supported by the extracted information.
Overall: In this case, both parts of the conclusion—"add GLP-1" and "continue Amlodipine"—are supported by the corresponding information. The model’s mention of "continuing Amlodipine" is a reasonable inference (most patients with T2DM have concurrent hypertension) but is not directly proven by the extracted information.
Therefore, the majority of the conclusion (the GLP-1 part) is fully supported by the information, while a small part of the conclusion (continuing Amlodipine) is a reasonable but unsupported inference. This fits the criteria for a "4-point" scenario.
Example 3 (3 points, Diagnosis)
1. Task Type:
Predict the patient’s diagnosis.
2. Model Reasoning Chain:
##Extraction:
Source: Outpatient Record [2179-05-10 10:30:00]
- Key Information: Patient complains of sore throat, cough, fever, and fatigue for 3 days.
- Analysis: Suggests symptoms of an upper respiratory tract infection (URTI).
Source: Physical Examination [2179-05-10 10:45:00]
- Key Information: Temperature 38.5°C, pharyngeal congestion, bilateral tonsillar enlargement (Grade II).
- Analysis: Supports acute tonsillitis or acute pharyngitis.
Source: Lab Results [2179-05-10 11:00:00]
- Key Information: White blood cell count 12.5×109/L (normal 4-10), Neutrophils 80% (normal 50-75%).
- Analysis: Both white blood cells and neutrophils are elevated, suggesting a bacterial infection.
##Reasoning:
Sore throat, fever, cough, temperature 38.5°C, pharyngeal congestion, tonsillar enlargement → Suggests URTI.
Elevated white blood cells and neutrophils → Suggests bacterial infection.
Combining the above information, the diagnosis is "acute upper respiratory tract infection with bacterial infection."
Given the symptoms and signs, acute bronchitis is also a possibility.
##Final Result:
Diagnosis: Acute upper respiratory tract infection with bacterial infection, acute bronchitis.
3. Score Explanation:
Score: 3 points
Reasoning:
The extracted information (sore throat, fever, elevated temperature, pharyngeal congestion, elevated white blood cells and neutrophils) fully supports the diagnosis of "acute upper respiratory tract infection with bacterial infection." This part of the conclusion is highly consistent with the information and is correct.
The diagnostic conclusion also includes "acute bronchitis," but the extracted information does not contain direct evidence to support this diagnosis. For example, there is no mention of worsening cough, productive cough, or abnormal findings on lung auscultation (such as rales or rhonchi).
While the majority of the conclusion ("acute upper respiratory tract infection with bacterial infection") is supported by the information, the diagnosis of "acute bronchitis" lacks a basis. According to the scoring criteria, "The extracted information only supports a small part of the conclusion, and the final result contains a large portion of content that is not adequately supported," therefore, the score is 3 points.
Example 4 (2 points, Medication Prescription)
1. Task Type:
Predict the patient’s medication plan.
II. Model Reasoning Chain:
##Extraction:
Source: Inpatient Record [2180-03-05 09:00:00]
- Key Information: Patient admitted due to palpitations and chest tightness. ECG shows atrial fibrillation.
- Analysis: The patient has a cardiac arrhythmia.
Source: Past Medical History [2178-01-15 14:00:00]
- Key Information: History of hypertension, long-term use of Irbesartan.
- Analysis: Blood pressure control needs to be continued.
Source: Lab Results [2180-03-05 10:30:00]
- Key Information: Normal kidney function, normal liver function.
- Analysis: Facilitates drug selection.
##Reasoning:
The patient has atrial fibrillation and needs heart rate control.
Long-term hypertension requires continued blood pressure management.
Considering all factors, Metoprolol is recommended.
##Final Result:
Prescription: Metoprolol.
3. Score Explanation:
Score: 2 points
Reasoning:
The extracted information—"palpitations, chest tightness, atrial fibrillation"—does indicate a cardiac arrhythmia, and Metoprolol is a commonly used first-line drug for treating atrial fibrillation. From a pharmacological perspective, this conclusion is related to the extracted information.
However, the extracted information does not provide sufficient evidence to support Metoprolol as the sole or preferred drug. The treatment plan for atrial fibrillation requires a comprehensive consideration of the patient’s heart rate control needs (type and rate of A-fib), stroke risk (CHA2DS2-VASc score), presence of comorbidities (like heart failure), and blood pressure levels. Based solely on "atrial fibrillation" and "hypertension," one cannot directly conclude that Metoprolol should be prescribed, as there are various other drug options (e.g., Diltiazem, Verapamil).
The conclusion "Metoprolol" lacks sufficient evidence to be fully supported. It might be a plausible treatment, but it is not the only or best choice that the extracted information can definitively determine. This aligns with the scoring criterion: "The extracted information is related to the final result, but it fails to support the conclusion."
Example 5 (1 point, Lab Test)
1. Task Type:
Predict the lab tests the patient should undergo.
2. Model Reasoning Chain:
##Extraction:
Source: Outpatient Record [2179-11-20 15:00:00]
- Key Information: Patient complains of abdominal pain and diarrhea.
- Analysis: Suggests a gastrointestinal issue.
Source: Physical Examination [2179-11-20 15:20:00]
- Key Information: Abdominal palpation shows no significant tenderness, bowel sounds are active.
- Analysis: Suggests no organic lesion or inflammation in the abdomen.
Source: Past Medication History [2179-11-19 18:00:00]
- Key Information: Recently took Aspirin.
- Analysis: No direct correlation with diarrhea.
##Reasoning:
Patient with abdominal pain and diarrhea → Gastrointestinal symptoms.
No tenderness on abdominal palpation → Rules out conditions like acute peritonitis.
Patient took Aspirin → May affect the gastrointestinal tract.
Based on the above information, an autoimmune disease is suspected.
##Final Result:
Recommended Tests: Antinuclear antibody (ANA) profile, Complement, Rheumatoid factor.
3. Score Explanation:
Score: 1 point
Reasoning:
The extracted information is "abdominal pain, diarrhea," "no tenderness on abdominal palpation," and "took Aspirin." This information points to common gastrointestinal symptoms and potential drug-related GI discomfort.
However, the final conclusion, "Recommended Tests: Antinuclear antibody profile, Complement, Rheumatoid factor," is used to diagnose autoimmune diseases (e.g., Systemic Lupus Erythematosus, Rheumatoid Arthritis).
There are no symptoms or signs in the extracted information (e.g., joint pain, rash, oral ulcers) that are related to autoimmune diseases. Linking "abdominal pain and diarrhea" to tests for autoimmune diseases lacks any reasonable basis.
The conclusion is severely irrelevant to the extracted information, and the reasoning process is highly unreliable. Therefore, the score is 1 point.
Prompt 2.
Reasoning Chain Synthesis Prompt for GPT-4o
# Patient EHR Context #
{context}
# Retrieved Medical Knowledge #
{medical_knowledge}
# Ground Truth #
{ground_truth}
# Task #
{task}
# Data Description
- # Patient EHR Context #: Contains all medical events and content from the patient’s hospitalization journey.
- # Retrieved Medical Knowledge #: Contains Patient EHR Context elements and their relationships potentially relevant to the Ground Truth entity.
- # Ground Truth #: Contains the correct answers for this task; your predictions must exactly match these.
- # Task #: Contains the specific task description; you need to complete the task based on the information in # Patient EHR Context #.
# Instructions
Please provide a logically rigorous medical reasoning process so that the # Ground Truth # can be derived from the content in # Patient EHR Context # and # Task #.
# Requirements
The reasoning process should include three stages: Extraction, Reasoning, and Final Results.
## Extraction
- In this stage, extract and identify each piece of "key information" from the # Patient EHR Context # according to the provided # Retrieved Medical Knowledge #.
- Don’t pay attention to information that you think is not helpful for the reasoning.
- Each step in the Extraction stage should follow the format below, **You need to specify the event name and time for each extracted information**:
**Event Name [Event Time]**: list the information extracted from the event and analyze the potential relationship between the key information and the ground truth.
## Reasoning
- Analyze the relationship between the context information and the item in # Ground Truth # in a very specific and professional manner, providing detailed reasoning steps.
- Your analysis should include the item in # Ground Truth # as much as possible. Items that cannot be inferred from the context can be omitted.
- Do not use the word "maybe", "possible" or "though" in the generated reasoning. You should do your best to find all the supporting information you can to ensure the correctness of your reasoning.
- The reasoning process should be concise and rigorous, and each step should explain the specific medical knowledge involved, making the reasoning process more credible.
- All reasoning must be based on the context information to infer the items in the ground truth and no reverse inference can be performed.
## Final Results - Provide the final result for the task. **Note that the final result should only contain the items contained in # Gound Truth # that have been correctly inferred in.
## Reasoning stage.**
- Each item in the ## Final Results should be contained in the # Gound Truth # with extactly same string.
# Important Notes!!!
- **For each piece of # Retrieved Medical Knowledge # that is relevant to completing the # Task #, locate the exact position of its first item within the # Patient EHR Context # and explicitly annotate it during the Extraction phase to ensure a more thorough analysis.**
- **During the ## Reasoning stage, remember to analyze very carefully how each item in # Ground Truth # is inferred**
- **Most importantly, integrate references to # Ground Truth # and # Retrieved Medical Knowledge # in an implicit manner. At any point in the reasoning process, do not use phrases such as “according to the medical knowledge above”, "as shown in ground truth" or any wording that reveals you are aware of the underlying medical knowledge or the ground truth.**
# Output Format
## Extraction
[YOUR OUTPUT]
## Reasoning
[YOUR OUTPUT]
## Final Results
[YOUR OUTPUT]